# The Green Peace Dividend: the Effects of Militarization on Emissions and the Green Transition

### v1

@@ 2.3 Discussion of empirical results
The empirical results robustly indicate that both emissions and emission intensity increases following a military spending shock equivalent to doubling the military spending share. In the main (NATO) sample, total emissions rise by 1.54%, while emission intensity rises by 1%. The data suggests that the rise in emissions is due to both a rise in emission intensity and a rise in output following a military spending shock. In the high-emission-intensity countries of the extended sample emissions and intensity also rise, both by 1%. However, low-intensity countries experience a drop in total emissions. I interpret this as evidence that emissions rise by 1-1.5% in response to a doubling of the military expenditure share, while emission intensity rises by 1%, with the qualification that this effect might be lower in low-energy- and emission-intensity economies.
Note that there are other potential margins of heterogeneity that I do not address in my method. As Ohanian (1997) pointed out, financing a war with debt can lead to a more pronounced rise in output than financing it with taxes. An upshot of this is that emissions could also rise more in the case of debt-financing. [^10] Unless financing methods affect emission intensity differentially. It is unclear how the defence shocks in the data were financed - this might impact the overall results.
An important difference might arise between temporary and permanent military spending shocks. My empirical strategy was aimed to uncover the effect of temporary shocks. However, as Barro (1981) pointed out, permanent government spending shocks might affect the economy differently compared to temporary shocks - in particular, their effect on output might be lower. [^11] A permanent military expenditure shock could materialize if, for example, the international security situation is perceived to worsen in the long run. In consequence, this would mean that the effect on total emissions in any given year of a permanent shock would be lower.
The conclusion that military buildups lead to a significantly higher volume of emissions and higher emission intensity is extremely concerning. In the extended sample, total emission intensity has been falling at an annualised rate of approximately 1.45% between 2000 and 2016, while total emissions have been falling at a rate of approximately 0.96% in the US. [^12] Total emissions have actually been rising in the extended sample in this period, at an annual rate of 1.88%. Hence, increasing the military spending share in the sample countries could erase several years of progress: for example, if the United States doubled its military expenditure share relative to 2023, from 3.4% to 6.8% [^13] SIPRI figures: https://milex.sipri.org/sipri, the results would indicate a 3.4-5.1% rise in total emissions.
To cross-check the empirical results and to understand the mechanisms through which a military spending shock affects emissions and the emission intensity, as well as the consequences of such a shock on the green transition, it’s necessary to use a more structured approach. The following section describes the economic model used for these purposes.
@@ 3 Model
This section describes the theoretical model. I use it to quantify the effect of a military buildup on greenhouse gas emissions, to understand how it impacts the green transition, as well as the mechanisms through which these effects work.
@@ 3.1 Setup
Where $C_{t}$ is aggregate consumption in period t, $L_{t}$ is aggregate labour supply, and $f_{it}$ is the household consumption of good variety i in period t. For simplicity, I assume log consumption utility, and a CES labour disutility function, where the Frisch elasticity of substitution is $\xi$. I assume a Cobb-Douglas consumption aggregator for simplicity.
@@
Where $E$ denotes total emissions, $\epsilon_{i}$ denotes the share of industry i’s scope 1 emissions171717”Scope 1 emissions are direct greenhouse (GHG) emissions that occur from sources that are controlled or owned by an organization (e.g., emissions associated with fuel combustion in boilers, furnaces, vehicles)” (US EPA) out of the total emissions of the economy, $\epsilon_{HH}$ denotes the share of household scope 1 emissions out of total emissions, and $\epsilon_{Govt}$ is the share of the government’s scope 1 emissions out of the total.
@@ 3.2 Calibration
Following vom Lehn and Winberry (2021), I add maintenance investment into the investment network. That is, I assume that each industry procures 12.5% of its total investment from itself - this can be interpreted as "maintenance investment". The reason for this adjustment is that the investment input-output matrix is singular [^20] Because investment sales are dominated by 4 main hubs, as described by vom Lehn and Winberry (2021)., and adding diagonal terms ensures its invertibility. Note that this adjustment is not necessary to calculate the steady-state results - the results are virtually the same without the adjustment in the steady-state calculations.
@@
I want to define the change in emissions as a function of the size of the military spending shock - that is, by how much does the military spending share of GDP change (as a multiple of the previous level). Define this spending multiplier as:
$M=\frac{S+E}{S}$
Where $S$ is the initial military spending share (a share of GDP), and $E$ is the additional spending share on the military. Now, the procurement multiplier - that is, by how much does government spending on weapon procurement change - is:
@@
$PM(M)=\frac{s_{P}\cdot S\cdot M+S_{P}-s_{P}\cdot S}{S_{P}}$
Where $S_{P}$ is the initial GDP share of government spending on weapon procurement, and $s_{P}$ is the share of additional total military expenditures spent on procurement. The second line follows from the definition of M. Similarly, the energy/fuel multiplier (defined analogously) is:
$EM(M)=\frac{s_{E}\cdot S\cdot M+S_{E}-s_{E}\cdot S}{S_{E}}$
@@
$\hat{G}_{it}=\begin{cases}1+\rho^{t-t_{0}}(PM(M)-1)&i\in\text{Weapon industries}\\ 1+\rho^{t-t_{0}}(EM(M)-1)&i\in\text{Energy/fuel industries}\\ 1&\text{ otherwise}\end{cases}$
@@
I consider three different calibrations - or scenarios - for $s_{P}$ and $s_{E}$. The baseline scenario has $s_{P}=0.3$ (30%) and $s_{E}=0.05$ (5%), which is higher than the 2017 share of spending on procurement and energy (around 20%, and around 2%, respectively (DoD (2016) and Crawford (2019)). This envisions a scenario where the government focuses on improving the striking power of the military by expanding its material base (number of weapons, ammunition, etc.) and by picking up the pace of military activity (e.g. through more frequent training or naval/air patrols) relative to the peacetime norm. Another reason why a relatively larger share of extra spending would be spent on fuel and weapons is that a war breaks out, and tanks, planes, etc. need to be used more intensively. This leads to a relatively higher rate of fuel use and a higher necessary replacement rate of weapons. [^25] Cottarelli and Virgadamo (2024) argue that EU countries will have to significantly increase the share of military spending on equipment, infrastructure, and operations to have well-equipped and prepared armed forces. In fact, they show that in the EU, 46.2% of the increase in defense expenditure since 2015 is due to expenditure on equipment and infrastructure.
The "personnel-oriented policy" calibration has $s_{P}=0.2$ and $s_{E}=0.02$ that is, the actual share of the 2017 budget spent on procurement and energy/fuel. This scenario assumes that it is going to be very hard for militaries to find enough manpower, and will need to spend a relatively large part of their extra budget on paying its extra personnel. The "material-oriented policy" calibration assumes $s_{P}=0.4$ and $s_{E}=0.10$ this scenario envisions a military that does not have difficulties with personnel and can focus its budget on military materials (fuel and weapons).
@@ 3.3.1 Permanent shock
Figure 10 shows the percentage change in total emissions, as a function of the size of the military spending share shock (as a multiple of the previous value, which was 3.3% in the US in 2017). The red line shows the baseline calibration, while the two blue lines show the "personnel-oriented" and the "material-oriented" calibrations. I also show two relevant scenarios with vertical dashed lines: the first shows the point at which the US government doubles its military spending share, while the second shows a scenario where the military spending share rises all the way to its Korean War peak (13.9% of GDP).
@@
Note: the figure shows how much total emissions change in the new steady state with the baseline calibration, relative to the old one, in response to a military spending shock (in percentage terms). The size of the spending shock, in multiples, is shown on the x-axis. The blue dashed lines show the results in the previously defined "personnel-oriented" calibration (lower line), and the "material-oriented" calibration (upper line). Two scenarios are marked with vertical dashed lines: the first is the point at which the military spending share is doubled, the second is where it is increased to its Korean War peak level.
@@
In response to a doubling of the military spending share, total emissions rise by approximately 3% in the new steady state under the baseline calibration. However, depending on the calibration of the shock, this can be both higher or lower. Under the "personnel-oriented" calibration, total emissions rise by only 1.18%, while under the "material-oriented" policy they rise by 6%. Also, note that under the extreme case of the military expenditure share rising to its Korean War peak, the baseline calibration predicts an almost 10% rise in total emissions, with an even wider "confidence interval" depending on the calibration.
As the relationship between the percentage change in emissions and the change in the military spending share is approximately linear (see figure 11), it is straightforward to translate this into a response of emissions to a percentage point rise in the military spending share. Using the fact that this share was 3.3% in 2017 in the US, we get a central estimate of a 0.9% rise in total emissions in response to a 1 percentage point rise in the military spending share, with a lower and upper bound of 0.36% and 1.81%, respectively.
Scenarios Emission intensity Total production Total effect
Doubling +2.24% +0.76% +3%
Korean War level +7.05% +2.55% +9.6%
Table 2: Change in total emissions, decomposed into change in emission intensity and change in total production. Rows show the two different militarization scenarios.
Table 2 shows that this effect is primarily due to a rise in the emission intensity of the economy, which is in line with the empirical evidence. Using the previous method, we get that a 1 percentage point rise in the military spending share raises emission intensity by between 0.22% and 1.5%, with a central estimate of 0.68% (divide 2.24% by 3.3%).
The model also predicts that real production rises and contributes to the increase in emissions, which is in line with the previous results. [^26] Changes in real GDP are calculated using a Laspeyres quantity index.
Scenarios Weapon procurement Energy/fuel use Total effect
Doubling -0.05% +3.05% +3%
Korean War level -0.16% +9.76% +9.6%
Table 3: Change in total emissions, decomposed into the effect of the weapon procurement and the energy/fuel demand shocks. Rows show the two different militarization scenarios.
@@
Table 3, however, shows that the primary reason behind the rise in total emissions is the energy/fuel demand shock. The weapon procurement shock decreases total emissions by a small amount - the explanation for this is that, in this model, the representative household reduces its demand for energy and fuel, which more than counteracts the effect of government demand for energy-intensive goods.
@@
Note that - as table 4 shows - one of the effects of the military spending shock is that, in the new steady state, energy-intensive industries, like primary metal producers, rubber and plastic firms (which all produce materials essential for weapon production), as well as the oil industry experience a significant boost. Services and agriculture lose the most. This could lead to an increase in the lobbying and political power of these energy-intensive and fossil fuel industries, which could make the green transition harder to accomplish.
Sullivan et al. (2022) argue that anti-regulatory lobbying helps those with abundant resources, while those lacking it can be penalised. The results here show that a military buildup can significantly boost the resources of polluting sectors. In addition, Eun et al. (2023) argue that firms with a larger share of green products out of total firm production conduct more intensive green lobbying. In case manufacturing firms (such as vehicle manufacturers) redirect production from green goods to military equipment, a militarization shock could translate into a drop in green lobbying efforts.
@@ 3.3.2 Temporary shock
Figure 11 shows the effect of the temporary shock on total emissions, in the two scenarios (doubling the military spending share, and bringing it to its Korean War peak level). The contemporary effect is 2.3% and 7.4%, respectively - somewhat less than the long-run effect of a military spending shock. Following that, emissions slowly return to their long-run trend. In case new spending is primarily material-oriented, this effect is much higher for longer, while if it’s personnel-oriented, it is lower, and returns to virtually zero after 4 years.
(a) Doubling
(b) Korean War level
@@
Figure 12: IRF of a temporary military spending shock’s effect on total emissions. The red line shows the baseline calibration, and the upper and lower blue dashed lines show the "material-oriented" and "personnel-oriented" calibrations, respectively. Figure a) shows a shock equivalent to doubling the military spending share, while figure b) shows a shock equivalent to increasing the share to its Korean War peak level.
Figure 12 shows that this effect is caused, primarily, by a rise in emission intensity. However, note that in case of a shock large enough to bring the military spending share to its Korean War peak level, the overall rise in production is responsible for almost half of the contemporary effect.
(a) Doubling
(b) Korean War level
@@
Figure 13: IRF of a temporary military spending shock’s effect on emission intensity. The red line shows the baseline calibration, and the upper and lower blue dashed lines show "material-oriented" and "personnel-oriented" calibrations, respectively. Figure a) shows a shock equivalent to doubling the military spending share, while figure b) shows a shock equivalent to increasing the share to its Korean War peak level.
@@
The mechanism is broadly similar to the one identified in the previous section, with a permanent shock. It is important, however, to note that the contemporary effect is also driven by a large investment shock, which mostly disappears as capital is reallocated to accommodate the new, defence-oriented economic structure. [^27] This temporary investment boom is partly responsible for the contemporary rise in the measured overall production, mentioned previously. In particular, figure 13 shows the evolution of the capital stock in two different industries: the "Other transport equipment" industry, which makes up a large fraction of the defence industry, and the utilities sector, in response to a shock that brings the military spending share to its Korean War peak level. Investment booms in the defence industries, with the capital stock peaking at more than 160% above its trend level 2 years after the shock. [^28] The IRFs show the evolution of the capital stock used in the next period - that is, the value of the IRF in period t reflects the movement in $K_{t+1}$. Meanwhile, the capital stock in the utilities sector stagnates in the first two periods, after which investment slumps, and the capital stock falls by 6% below trend, 7 years after the shock. In general, investments in the defence industries crowd out investments in other sectors of the economy. [^29] Crowd-out of investment in utilities happens even though the military spending shock has a positive effect on it. As lump-sum taxes are used to fund new expenditures, the household reduces its spending on energy significantly. In addition, some other industries which are negatively affected also reduce their energy demand.
@@
The reason why it is particularly important to emphasize the slump in investment in the utilities sector is that new investments in this industry usually go to renewable projects - hence, a crowd-out of investment from this sector could retard the green transition by slowing down investment in green energy sources. In addition, if the military spending shock also crowds out investment into green research and development - something that this model doesn’t capture directly -, that could further slow and reduce the chances of a green transition (see e.g. Acemoglu et al. (2012, 2024)).
@@
Energy-intensive industries, as well as fossil-fuel industries, experience a significant boost in size and activity, while investment and production in other sectors are crowded out. This can lead to a fall in green investment and green innovation, which could slow the green transition. It can also lead to a rise in the political power of energy-intensive and fossil-fuel industries, which could reduce the chance of passing and implementing crucial climate policies.
4 The Green Peace Dividend

### v3

@@ 2.3 Discussion of empirical results
The empirical results robustly indicate that both emissions and emission intensity increases following a military spending shock equivalent to doubling the military spending share. In the main (NATO) sample, total emissions rise by 1.94%, while emission intensity rises by 0.9%. The data suggests that the rise in emissions is due to both a rise in emission intensity and a rise in output following a military spending shock. In the high-emission-intensity countries of the extended sample emissions and intensity also rise, the former by 0.9%, the latter by more than 1%. However, low-intensity countries experience a drop in total emissions. I interpret this as evidence that emissions rise by 0.9-2% in response to a doubling of the military expenditure share, while emission intensity rises by 1%, with the qualification that this effect might be lower in low energy and emission intensity economies.
The results also suggest that a military spending shock reduces green patenting by 10-25%, either due to crowding out innovation in this field or as a result of a general negative effect on patenting.
Note that there are other potential margins of heterogeneity that I do not address in my method. As Ohanian (1997) pointed out, financing wars with debt can lead to a more pronounced rise in output than financing it with taxes. An upshot of this is that emissions could also rise more in the case of debt-financing. [^14] Unless financing methods affect emission intensity differentially. It is unclear how the defence shocks in the data were financed - this might impact the overall results.
An important difference might arise between temporary and permanent military spending shocks. My empirical strategy was aimed at uncovering the effect of temporary shocks. However, as Barro (1981) pointed out, permanent government spending shocks might affect the economy differently compared to temporary shocks - in particular, their effect on output might be lower. [^15] A permanent military expenditure shock could materialize if, for example, the international security situation is perceived to worsen in the long run. In consequence, this would mean that the effect on total emissions in any given year of a permanent shock would be lower.
To cross-check the empirical results and to better understand the mechanisms through which a military spending shock affects emissions and the emission intensity, as well as the consequences of such a shock on the green transition, I use a dynamic production network model.
@@ 3 Model
This section describes the theoretical model. I use it to provide additional evidence on the effect of a military buildup on greenhouse gas emissions, to understand what other channels it impacts the green transition through, as well as the mechanisms through which these effects work.
@@ 3.1 Setup
I do not discuss innovation in the context of my model. The primary reason is that that would make it significantly more complicated, as it would require modelling the market power of the industries and taking a stance on the relevant innovation process, which is beyond this paper’s scope.
@@
Where $C_{t}$ is aggregate consumption in period t, $L_{t}$ is aggregate labour supply, and $f_{it}$ is the household consumption of good variety i in period t. For simplicity, I assume log consumption utility, and an additively separable labour disutility function, where the Frisch elasticity of substitution is $\xi$. I assume a Cobb-Douglas consumption aggregator for simplicity.
@@
Where $E$ denotes total emissions, $\epsilon_{i}$ denotes the share of industry i’s scope 1 emissions191919”Scope 1 emissions are direct greenhouse (GHG) emissions that occur from sources that are controlled or owned by an organization (e.g., emissions associated with fuel combustion in boilers, furnaces, vehicles)” (US EPA) out of the total emissions of the economy, $\epsilon_{HH}$ denotes the share of household scope 1 emissions out of total emissions, and $\epsilon_{Govt}$ is the share of the government’s scope 1 emissions out of the total.
@@ 3.2 Calibration
Following vom Lehn and Winberry (2021), I add maintenance investment into the investment network. That is, I assume that each industry procures 12.5% of its total investment from itself - this can be interpreted as "maintenance investment". The reason for this adjustment is that the investment input-output matrix is singular [^22] Because investment sales are dominated by 4 main hubs, as described by vom Lehn and Winberry (2021)., and adding diagonal terms ensures its invertibility. Note that this adjustment is not necessary to calculate the steady-state results - the results are virtually the same without the adjustment in the steady-state calculations. However, it is necessary for calculating the dynamic effects of a temporary shock.
@@
I want to define the change in emissions as a function of the size of the military spending shock - that is, by how much does the military spending share of GDP change (as a multiple of the previous level). Now, the procurement multiplier - that is, by how much does government spending on weapon procurement change - is:
@@
Where $S_{P}$ is the initial GDP share of government spending on weapon procurement, $s_{P}$ is the share of additional total military expenditures spent on procurement, and E is the rise in the total military expenditure share, in percentage points. Similarly, the energy/fuel multiplier (defined analogously) is:
$EM=\frac{S_{E}+s_{E}E}{S_{E}}$
@@
$\hat{G}_{it}=\begin{cases}1+\rho^{t-t_{0}}(PM-1)&i\in\text{Weaponindustries}\\ 1+\rho^{t-t_{0}}(EM-1)&i\in\text{Energy/fuel industries}\\ 1&\text{ otherwise}\end{cases}$
@@
Since there is uncertainty regarding the share of potential future defence spending shocks going to equipment and energy, I consider three different calibrations - or scenarios - for $s_{P}$ and $s_{E}$. The baseline scenario has $s_{P}=0.3$ (30%) and $s_{E}=0.05$ (5%), which is higher than the 2017 share of spending on procurement and energy (around 20%, and around 2%, respectively (DoD (2016) and Crawford (2019)). This envisions a scenario where the government focuses on improving the striking power of the military by expanding its material base (number of weapons, ammunition, etc.) and by picking up the pace of military activity (e.g. through more frequent training or naval/air patrols) relative to the peacetime norm. Another reason why a relatively larger share of extra spending would be spent on fuel and weapons is that a war breaks out, and tanks, planes, etc. need to be used more intensively. This leads to a relatively higher rate of fuel use and a higher replacement rate of weapons. [^27] Cottarelli and Virgadamo (2024) argue that EU countries will have to significantly increase the share of military spending on equipment, infrastructure, and operations to have well-equipped and prepared armed forces. In fact, they show that in the EU, 46.2% of the increase in defense expenditure since 2015 is due to expenditure on equipment and infrastructure.
The "personnel-oriented policy" calibration has $s_{P}=0.2$ and $s_{E}=0.02$ that is, the actual share of the 2017 budget spent on procurement and energy/fuel. This scenario assumes that it is going to be very hard for militaries to find enough manpower, and will need to spend a relatively large part of their extra budget on paying its extra personnel. The "material-oriented policy" calibration assumes $s_{P}=0.4$ and $s_{E}=0.1$ this scenario envisions a military that does not have difficulties with personnel and can focus its budget on military materials (fuel and weapons).
@@ 3.3.1 Permanent shock
Calibration Total emissions Emission intensity Real GDP
Material-oriented +1.81% +1.5% +0.31%
Baseline +0.9% +0.68% +0.22%
Personnel-oriented +0.36% +0.22% +0.14%
Table 2: Change in total emissions, emission intensity, and real GDP in response to a permanent, 1 p.p. rise in the military spending share under the different shock calibrations.
Table 2 shows the response of GHG emissions and emission intensity in the new steady-state, following a permanent 1 p.p. shock to the military spending share. Total emissions rise by between 0.36% and 1.81% (depending on the shock composition), while emission intensity rises by 0.22-1.5%. Four important points should be made here. First, the rise in emission intensity makes up 75-80% of the total rise in emissions, which means that, in the model, the total effect is not simply due to a positive spending multiplier. Second, real GDP marginally increases under all calibrations, the fundamental reason being that labour supply increases following a rise in the lump-sum tax, due to the income effect induced by the higher tax. [^28] Changes in real GDP are calculated using a Laspeyres quantity index. Vom Lehn and Winberry (2021) use a Törnqvist index in their simulations.
Third, there is a significant overlap between the baseline empirical estimate of an effect of 0.9-2% and the model results, although the interval is shifted somewhat downwards. Fourth, the results are sensitive to the shock calibration - there is a difference of 1.45% between the effect in a personnel-oriented and a material-oriented scenario.
To underscore this uncertainty, figure 13 shows the percentage change in total emissions, as a function of the size of the military spending share shock (in percentage points). The red line shows the baseline calibration, while the two blue lines show the "personnel-oriented" and the "material-oriented" calibrations. I also show two relevant scenarios with vertical dashed lines: the first shows the point at which the US government doubles its military spending share (from a baseline of 3.3% in 2017), while the second shows a scenario where the military spending share rises to its Korean War peak (13.9% of GDP).
@@
Note: the figure shows how much total emissions change in the new steady state with the baseline calibration, relative to the old one, in response to a military spending shock (in percentage terms). The size of the spending shock, in percentage points, is shown on the x-axis. The blue dashed lines show the results in the previously defined "personnel-oriented" calibration (lower line), and the "material-oriented" calibration (upper line). Two scenarios are marked with vertical dashed lines: the first is the point at which the military spending share is doubled, and the second is where it is increased to its Korean War peak level.
@@
In response to a doubling of the military spending share, total emissions rise by approximately 3% in the new steady state under the baseline calibration. However, depending on the calibration of the shock, this can be both higher or lower. Under the "personnel-oriented" calibration, total emissions rise by only 1.18%, while under the "material-oriented" policy they rise by 6%. Also, note that under the extreme case of the military expenditure share rising to its Korean War peak, the baseline calibration predicts an almost 10% rise in total emissions, with an even wider "confidence interval" - under the material-oriented calibration, emissions rise by 20%.
From a military point of view (see e.g. Cottarelli and Virgadamo (2024)), the higher the share of expenditures spent on equipment and training, the more efficient and powerful the armed forces are. Hence, the material-oriented scenario better describes a scenario in which, for example, European countries can efficiently invest and improve their militaries. However, in the case of political inefficiencies in constructing and executing defence budgets, an economy can be closer to the personnel-oriented scenario, with small effects on GHG emissions.
Shocks Emission intensity Real GDP Total effect
Weapon procurement -0.2% +0.18% -0.02%
Energy/fuel use +0.88% +0.04% +0.92%
Table 3: Change in total emissions, emission intensity, and real GDP due to the weapon procurement and the energy/fuel use shocks, separately. A military expenditure shock of 1 p.p. under the baseline shock calibration is assumed. The results in either row are calculated by switching off the other shock.
Table 3 decomposes the effect of a 1 p.p. military expenditure share shock into the effect of the two components, the weapon procurement shock and the energy/fuel use shock. I do this by shutting down each shock in turn and looking at the effect of the remaining shock on emission intensity and output.
@@
The results show that the primary reason behind the rise in total emissions is the energy/fuel demand shock. The weapon procurement shock decreases total emissions by a small amount - the explanation for this is that, in this model, the representative household reduces its demand for energy and fuel, which more than counteracts the effect of government demand for energy-intensive goods. Real GDP increases for both shocks, which is simply the outcome of the income effect - note that the effect of the weapon procurement shock on real GDP is higher because it makes up 30% of the total expenditure shock under the baseline calibration, while the energy demand shock is just 5% of the total.
@@
Note that - as table 4 shows - one of the effects of the military spending shock is that, in the new steady state, energy-intensive industries, like primary metal producers, rubber and plastic firms (which all produce materials essential for weapon production), as well as the oil industry experience a significant boost. Services and agriculture lose the most. This could lead to an increase in the lobbying and political power of these energy-intensive and fossil fuel industries, which could make the green transition harder to accomplish (see, e.g., Sullivan et al. (2022) or Pacca et al. (2021)).
@@ 3.3.2 Temporary shock
Figure 14 shows the effect of the temporary shock of 1 percentage point on total emissions. The contemporary effect in the baseline shock calibration is approximately 0.7% - somewhat less than the long-run effect of a military spending shock. Following that, emissions slowly return to their long-run trend. In case new spending is primarily material-oriented, this effect is much higher (almost 1.5% contemporaneously) for longer, while if it’s personnel-oriented, it is lower, and returns to virtually zero after 4 years.
@@
Figure 14: IRF of a 1 p.p. temporary military spending share shock’s effect on total emissions. The red line shows the baseline calibration, and the upper and lower blue dashed lines show the "material-oriented" and "personnel-oriented" calibrations, respectively.
Figure 12 shows that this effect is caused, primarily, by a rise in emission intensity: emission intensity rises by more than 0.4% contemporaneously for the baseline calibration, about 57% of the total effect.
@@
Figure 15: IRF of a temporary military spending shock’s effect on emission intensity. The red line shows the baseline calibration, and the upper and lower blue dashed lines show "material-oriented" and "personnel-oriented" calibrations, respectively.
@@
The mechanism is broadly similar to the one identified in the previous section, with a permanent shock. It is important, however, to note that the contemporary effect is also driven by a large investment shock, which mostly disappears as capital is reallocated to accommodate the new, defence-oriented economic structure. [^29] This temporary investment boom is partly responsible for the contemporary rise in the measured overall production, mentioned previously. In particular, figure 13 shows the evolution of the capital stock in two different industries: the "Other transport equipment" industry, which makes up a large fraction of the defence industry, and the power sector, in response to a shock that brings the military spending share to its Korean War peak level. Investment booms in the defence industries, with the capital stock peaking at more than 160% above its trend level 2 years after the shock. [^30] The IRFs show the evolution of the capital stock used in the next period - that is, the value of the IRF in period t reflects the movement in $K_{t+1}$. Meanwhile, the capital stock in the utilities sector stagnates in the first two periods, after which investment slumps, and the capital stock falls by 6% below trend, 7 years after the shock. In general, investments in the defence industries crowd out investments in other sectors of the economy. [^31] Crowd-out of investment in utilities happens even though the military spending shock has a positive effect on it. As lump-sum taxes are used to fund new expenditures, the household reduces its spending on energy significantly. In addition, some other industries which are negatively affected also reduce their energy demand.
@@
The reason why it is particularly important to emphasize the slump in investment in the utilities sector is that new investments in this industry usually go to renewable projects (see IEA (2023): investment in renewables was more than half of all investments in the power sector globally, while investment in fossil fuels was less than 10%) - hence, a crowd-out of investment from this sector could retard the green transition by slowing down investment in green energy sources. Note, though, that investment in the power sector increases with a permanent shock in the new steady state. This suggests that whether investment in renewables falls or not might depend on the persistence of the rise in military spending. However, in case investment in the power sector increases in a new steady state, it is unclear whether the extra investment will flow to renewables, electrical grids, or storage technologies, or whether it will flow to dirty plants.
@@
Energy-intensive industries, as well as fossil-fuel industries, experience a significant boost in size and activity, while investment and production in other sectors are crowded out. This can lead to a fall in the investment in renewables, which could slow the green transition.
4 The Green Peace Dividend and policy implications